# Statistical Taylor Expansion111

### v1

@@ 8.5 Using the Indexed Sine Functions for Sin/Cos Signals
figure The result error deviation and uncertainty mean of Sin/Cos signals vs. FFT order using the indexed sine functions for forward, reverse and roundtrip FFT transformations, as shown in the legend. The y-axis on the left is for error deviation, while the y-axis on the right is for uncertainty mean.
@@
figure The histograms of the normalized errors of Sin/Cos signals using the indexed sine functions for forward, reverse and roundtrip FFT transformations, as shown in the legend. The FFT order is $18$.
Using the indexed sine functions, for the waveform of $\sin(2\pi k3/2^{6})$ in which $k$ is the index time:
@@
Figure 8.5 shows both the result uncertainty means and the error deviations vs. FFT order of FFT transformations of Sin/Cos signals using the indexed sine functions. As expected, the uncertainties grow much faster with increasing FFT order for the forward transformation. The error deviations are a constant about $0.65$ for the forward transformation, while they increase with increasing FFT order from $0.25$ until reach a constant about $0.85$ when the FFT order is $12$ or more for the reverse and roundtrip transformation.
@@ 8.6 Using the Library Sine Functions for Sin/Cos Signals
figure The FFT waveform of $\sin(j3/2^{6}\pi)$ using the indexed sine functions after the reverse transformation calculated by variance arithmetic, with the uncertainties and the value errors shown in the legend. Also included are the corresponding result value errors using SciPy. The x-axis is for index time. The y-axis is for uncertainties or absolute value errors.
@@
figure The result error deviation and uncertainty mean of Sin/Cos signals vs. FFT order using the library sine functions for forward, reverse and roundtrip FFT transformations, as shown in the legend. The y-axis on the left is for error deviation, while the y-axis on the right is for uncertainty mean.
@@
figure The result error deviation and uncertainty mean of Sin/Cos signals vs. FFT order using the library sine functions for forward, reverse and roundtrip FFT transformations, as shown in the legend. The y-axis on the left is for error deviation, while the y-axis on the right is for uncertainty mean. $10^{-16}$ input noises are added to the Sin/Cos signals.
@@
Because the least significant values are the only source of input uncertainties for variance arithmetic, the result uncertainties using either sine functions are almost identical. The library sine functions have numerical calculation errors which is not specified by the input uncertainties of variance arithmetic. The question is whether variance arithmetic can track these numerical errors effectively or not.
Using the library sine functions, for the waveform of $\sin(2\pi k3/2^{6})$ in which $k$ is the index time:
@@
  • Figure 8.5 and Figure 8.6 confirm that the result uncertainties are identical to all FFT transformations, respectively.
Variance arithmetic reveals the reason why to add small noises to the numerically generated input data, which is already a common practice [13]. It further reveals that the amount of noise needed is to achieve proper coverage.
In Figure 8.6, the value errors tend to increase with the index time, which is due to the periodic increase of the numerical errors of $\sin(x)$ with $x$ as shown in Figure 8.4. When the signal frequency increases, the increase of the value errors with the index frequency in the reverse transformation becomes stronger, as shown in Figure 8.6, 8.6, and 8.6. In fact, Figure 8.6 suggests that the periodic numerical errors in Figure 8.4 resonate with the periodicity of the input $Sin$ signal. In contrast, Figure 8.5 shows no indication for such increase. It shows that the library numerical errors may have surprisingly large effect on the numerical results.
@@ 8.7 Linear Signal
figure The result error deviation and uncertainty mean of Linear signal vs. FFT order using the indexed sine functions for forward, reverse and roundtrip FFT transformations, as shown in the legend. The y-axis on the left is for error deviation, while the y-axis on the right is for uncertainty mean.
@@
figure The histograms of the normalized errors of Linear signal using the indexed sine functions for forward, reverse and roundtrip FFT transformations, as shown in the legend.
@@
figure The result error deviation and uncertainty mean of Linear signal vs. FFT order using the library sine functions for forward, reverse and roundtrip FFT transformations, as shown in the legend. The y-axis on the left is for error deviation, while the y-axis on the right is for uncertainty mean.
@@
figure The histograms of the normalized errors of Linear signal using the library sine functions for forward, reverse and roundtrip FFT transformations, as shown in the legend.
@@
  • Figure 8.7 shows that the result uncertainty for each FFT transformation has much larger increase with the increasing FFT order than its counter part in Figure 8.5. The much faster uncertainty increase of the forward transformation is attributed to the non-zero input uncertainties. Figure 8.7 shows that proper coverage can be achieved for all FFT transformations for all FFT orders.
@@
  • Figure 8.7 shows that proper coverage cannot be achieved for all FFT transformations for all FFT orders, because the value errors outpace the uncertainties with increasing FFT order, to result in worse error deviations for increasing FFT order.
  • Figure 8.7 shows that the reverse histogram contains additional peaks to its counterpart in Figure 8.7. The additional peaks extend quite beyond the range of $[-3,+3]$, revealing the reason why variance arithmetic can not cover these value errors. Such large numerical errors are expected from Figure 8.4.
@@ 8.8 Ideal Coverage
figure The histograms of the normalized errors of Linear signal with $10^{-3}$ input noise using the library sine functions for forward, reverse and roundtrip FFT transformations, as shown in the legend.
@@
figure The result error deviation and uncertainty mean of Linear signal with $10^{-3}$ input noise using the library sine functions vs. FFT order for forward, reverse and roundtrip FFT transformations. The y-axis on the left is for error deviation, while the y-axis on the right is for uncertainty mean.
@@
  • Figure 8.8 shows the corresponding histogram. As expected, the normalized errors for the forward and the reverse transformations are Normal distributed, even when the input noise is not Gaussian. The normalized errors for the roundtrip transformations are Delta distributed at 00, meaning the input uncertainties are perfectly recovered.
@@
Also, the result uncertainty means for both the forward and the reverse transformations are linear to the input uncertainties, respectively, which is expected because FFT transformations are linear.
@@
In contrast, for Linear signals using indexed sine functions, as shown by Figure 8.8 and 8.8 for the forward and reverse transformations, respectively, the ideal region is much larger, and proper coverage is achieved in other regions. Forward FFT transformation still shows larger region of ideal coverage than that of Reverse FFT transformation.
For Sin/Cos using either indexed sine functions or library sine functions, the ideal region is achieved when the added noise is large enough to cover the effect of rounding errors, almost independent of FFT orders. In both cases, forward transformation needs $10^{-15}$ added noise, while reverse transformation needs $10^{-12}$ added noise. The difference is that using indexed sine functions, in the proper coverage region, error deviations deviate from $1$ only slightly.
@@ 8.9 Summary
Compared to its counterpart continuous Fourier Transformation (FT), the discrete Fourier transformation (DFT) has large modeling error due to its implied assumption. This modeling error has not been addressed seriously.
The library sine functions using conventional floating-point arithmetic have been shown to contain numerical errors as large as equivalently $10^{-3}$ of input precision for FFT transformations. The library $\sin(x)$ errors increase periodically with $x$, causing noticeable increase of the result errors with increasing index frequency in the reverse transformation, even in resonant fashion. The dependency of the result numerical errors on the amount of calculation and input data means that a small-scale test cannot properly qualify the result of a large-scale calculation. The effect of numerical errors inside math library has not been addressed seriously.
@@ 9 Regressive Generation of Sin and Cos
Starting from Formula (9.1), Formula (9.2) and Formula (9.3) can be used recursively to calculate the $\sin$ library function.
@@
The number of regressions is defined as the order. For each order $n$, $2^{n}$ $\sin(\frac{\pi}{2}\frac{i}{2^{n}})$ and $\cos(\frac{\pi}{2}\frac{i}{2^{n}})$ values are obtained, so that the result is more statistically stable with increasing $n$.
@@ 10.2 Summary of Variance Arithmetic
The presence of ideal coverage is the necessary condition for a numerical algorithm in variance arithmetic to be correct. The ideal coverage also defines the ideal applicable range for the algorithm. In ideal coverage, the calculated uncertainty equals the error deviation, and the result normalized errors is either Normal distributed or Delta distributed depending on the context. In variance arithmetic, the best-possible precision to achieve ideal coverage is $10^{-15}$, which is good enough for most applications.
Variance arithmetic can only provide proper coverage for floating-point rounding errors, with the empirical error deviations in the approximate range of $[1/3,3]$. Table 10.2 shows the measured error deviations in different contexts.
@@ 10.3 Improvement Needed
  • What is the uncertainty upper bounds for each problem? Specifically, how to explain quantitatively the measured upper bounds in Figure 3.4 and 3.4.
  • How to analytically quantify the reliability requirements for variance arithmetic?
@@
  • In variance arithmetic, a statistical Taylor expansion is invalidated as not reliable if the calculated variance has value less than $1/\sigma$-fold of its uncertainty. Such cutoff is somewhat arbitrary, and more mathematically rigid approach is needed.
10.4 New Numerical Approaches
Traditional numerical approaches focus primarily on the result values, and they need to be reexamined or even reinvented in variance arithmetic. Most traditional numerical algorithms try to choose one of the best paths, while variance arithmetic rejects conceptually all path-dependent calculations.
Variance arithmetic generally reveals worse results than those of conventional floating-point arithmetic, such as very quick loss of precision in matrix inversion, especially for matrix of large size. How to reconcile variance arithmetic and traditional numerical approaches could be a big challenge.
Variance arithmetic no longer has dependency problem, and it can reject ill-formed problems. By using stability truncation, variance arithmetic no longer needs theoretical reminder estimator so that it is simpler for truncation problems. Variance arithmetic seems conceptually easier than traditional numerical approaches.
The calculation of the result uncertainty bias and uncertainty contains a lot of sums of constant terms such as Formula (2.16) and (2.19), so that it is a excellent candidate for parallel processing. Because variance arithmetic has neither execution freedoms nor the associated dependency problems, it can be implemented either in software using GPU or directly in hardware.
In variance arithmetic, it is generally an order-of-magnitude more complicated to calculate the result uncertainties than the result values, such as Formula (6.19) for Taylor expansion of matrix inversion. However, modern software programs for analytic calculation such as SymPy can be a great help, whose usage is demonstrated in the corresponding open source project. Perhaps it is time to move from numerical programming to analytic programming.
@@ 10.9 Variable $\sigma$ Variance Arithmetic
Variance arithmetic rejects a calculation if its distributional pole or distributional zero is within the input ranges. But there are many use cases in which such intrusion cannot be avoided, so variance arithmetic of smaller $\sigma$ may be needed. It is even possible in another implementation of variance arithmetic, $\sigma$ is part of a variance representation, so that a calculation closer to a distributional zero or a distributional pole can proceed but at the expense of reducing $\sigma$. The result $\sigma$ for a calculation of imprecise inputs each with different $\sigma$ is found statistically through the corresponding $\epsilon$. Such approaches need theoretical foundation in both mathematics and statistics.

### v2

@@ 8.5 Using the Indexed Sine Functions for Sin/Cos Signals
figure The result error deviation and uncertainty mean of Sin/Cos signals vs. FFT order using the indexed sine functions for forward, reverse and roundtrip FFT, as shown in the legend. The y-axis on the left is for error deviation, while the y-axis on the right is for uncertainty mean.
@@
figure The histograms of the normalized errors of Sin/Cos signals using the indexed sine functions for forward, reverse and roundtrip FFT, as shown in the legend. The FFT order is $18$.
Using the indexed sine functions, for the waveform of $\sin(2\pi j3/2^{6})$ in which $j$ is the index time:
@@
Figure 8.5 shows both the result uncertainty means and the error deviations vs. FFT order of DTF of Sin/Cos signals using the indexed sine functions. As expected, the uncertainties grow much faster with increasing FFT order for the forward transformation. The error deviations are a constant about $0.65$ for the forward transformation, while they increase with increasing FFT order from $0.25$ until reach a constant about $0.85$ when the FFT order is $12$ or more for the reverse and roundtrip transformation.
@@ 8.6 Using the Library Sine Functions for Sin/Cos Signals
figure The FFT waveform of $\sin(j3/2^{6}\pi)$ using the library sine functions after the reverse transformation calculated by variance arithmetic, with the uncertainties and the value errors shown in the legend. Also included are the corresponding result value errors using SciPy. The x-axis is for index time. The y-axis is for uncertainties or absolute value errors.
@@
figure The result error deviation and uncertainty mean of Sin/Cos signals vs. FFT order using the library sine functions for forward, reverse and roundtrip FFT, as shown in the legend. The y-axis on the left is for error deviation, while the y-axis on the right is for uncertainty mean.
@@
figure The result error deviation and uncertainty mean of Sin/Cos signals vs. FFT order using the library sine functions for forward, reverse and roundtrip FFT, as shown in the legend. The y-axis on the left is for error deviation, while the y-axis on the right is for uncertainty mean. $10^{-16}$ input noises are added to the Sin/Cos signals.
@@
Because the least significant values are the only source of input uncertainties for variance arithmetic, the result uncertainties using either indexed or library sine functions are almost identical. The library sine functions have numerical calculation errors which are not specified by the input uncertainties of variance arithmetic. The question is whether variance arithmetic can track these numerical errors effectively or not.
Using the library sine functions, for the waveform of $\sin(2\pi j3/2^{6})$ in which $j$ is the index time:
@@
  • Figure 8.5 and Figure 8.6 confirm that the result uncertainties are identical to all FFT, respectively.
Variance arithmetic reveals why to add small noises to the numerically generated input data, which is already a common practice [13]. It further reveals that the amount of noise needed is to achieve proper coverage.
In Figure 8.6, the value errors tend to increase with the index time, which is due to the periodic increase of the numerical errors of $\sin(x)$ with $x$ as shown in Figure 8.4. When the signal frequency increases, the increase of the value errors with the index frequency in the reverse transformation becomes stronger, as shown in Figure 8.6, 8.6, and 8.6. In fact, Figure 8.6 suggests that the periodic numerical errors in Figure 8.4 resonate with the periodicity of the input $Sin$ signal. In contrast, Figure 8.5 shows no sign for such increase. It shows that the library numerical errors may have surprisingly large effect on the numerical results.
@@ 8.7 Linear Signal
figure The result error deviation and uncertainty mean of Linear signal vs. FFT order using the indexed sine functions for forward, reverse and roundtrip FFT, as shown in the legend. The y-axis on the left is for error deviation, while the y-axis on the right is for uncertainty mean.
@@
figure The histograms of the normalized errors of Linear signal using the indexed sine functions for forward, reverse and roundtrip FFT, as shown in the legend.
@@
figure The result error deviation and uncertainty mean of Linear signal vs. FFT order using the library sine functions for forward, reverse and roundtrip FFT, as shown in the legend. The y-axis on the left is for error deviation, while the y-axis on the right is for uncertainty mean.
@@
figure The histograms of the normalized errors of Linear signal using the library sine functions for forward, reverse and roundtrip FFT, as shown in the legend.
@@
  • Figure 8.7 shows that the result uncertainty for each FFT transformation has much larger increase with the increasing FFT order than its counterpart in Figure 8.5. The much faster uncertainty increase of the forward transformation is attributed to the non-zero input uncertainties. Figure 8.7 shows that proper coverage can be achieved for all FFT for all FFT orders.
@@
  • Figure 8.7 shows that proper coverage cannot be achieved for all FFT for all FFT orders, because the value errors outpace the uncertainties with increasing FFT order, to result in worse error deviations for increasing FFT order.
  • Figure 8.7 shows that the reverse histogram contains additional peaks to its counterpart in Figure 8.7. The additional peaks extend quite beyond the range of $[-3,+3]$, revealing why variance arithmetic cannot cover these value errors. Such large numerical errors are expected from Figure 8.4.
@@ 8.8 Ideal Coverage
figure The histograms of the normalized errors of Linear signal with $10^{-3}$ input noise using the library sine functions for forward, reverse and roundtrip FFT, as shown in the legend.
@@
figure The result error deviation and uncertainty mean of Linear signal with $10^{-3}$ input noise using the library sine functions vs. FFT order for forward, reverse and roundtrip FFT. The y-axis on the left is for error deviation, while the y-axis on the right is for uncertainty mean.
@@
  • Figure 8.8 shows the corresponding histogram. As expected, the normalized errors for the forward and the reverse transformations are Normal distributed. The normalized errors for the roundtrip transformations are Delta distributed at 00, meaning the input uncertainties are perfectly recovered.
@@
Also, the result uncertainty means for both the forward and the reverse transformations are linear to the input uncertainties, respectively, which is expected because FFT is linear.
@@
In contrast, for Linear signals using indexed sine functions, as shown by Figure 8.8 and 8.8 for the forward and reverse transformations, respectively, the ideal region is much larger, and proper coverage is achieved in other regions. Forward FFT transformation shows larger region of ideal coverage than that of Reverse FFT transformation.
As an comparison, because sin/cos has less numerical calculation errors, for Sin/Cos using either indexed sine functions or library sine functions, the ideal region is achieved when the added noise is large enough to cover the effect of rounding errors, almost independent of FFT orders. In both cases, forward transformation needs $10^{-15}$ added noise, while reverse transformation needs $10^{-12}$ added noise. The difference is that using indexed sine functions, in the proper coverage region, error deviations deviate from $1$ only slightly.
@@ 8.9 Summary
The library sine functions using conventional floating-point arithmetic have been shown to contain numerical errors as large as equivalently $10^{-3}$ of input precision for FFT. The library $\sin(x)$ errors increase periodically with $x$, causing noticeable increase of the result errors with increasing index frequency in the reverse transformation, even in resonant fashion. The dependency of the result numerical errors on the amount of calculation and input data means that a small-scale test cannot properly qualify the result of a large-scale calculation. The effect of numerical errors inside math library has not been addressed seriously.
@@ 9 Regressive Generation of Sin and Cos
[…]
Formula (9.2) and Formula (9.3) calculate $\sin(j\frac{\pi}{2n}),\cos(j\frac{\pi}{2n}),j=0\dots N$ repressively starting from Formula (9.1), .
@@
The number of regressions is defined as the order. For each order $n$, $2^{n}$ $\sin(j\frac{\pi}{2n})$ and $\cos(j\frac{\pi}{2n})$ values are obtained, so that the result is more statistically stable with increasing $n$.
@@ 10.2 Summary of Variance Arithmetic
The presence of ideal coverage is the necessary condition for a numerical algorithm in variance arithmetic to be correct. The ideal coverage also defines the ideal applicable range for the algorithm. In ideal coverage, the calculated uncertainty equals the value error deviation, and the result normalized errors is either Normal distributed or Delta distributed depending on the context. In variance arithmetic, the best-possible precision to achieve ideal coverage is $10^{-15}$, which is good enough for most applications.
Variance arithmetic can only provide proper coverage for floating-point rounding errors, with the empirical error deviations in the approximate range of $[1/5,5]$. Table 10.2 shows the measured error deviations in different contexts.
@@ 10.3 Improvement Needed
  • What is the uncertainty upper bounds for each problem? Specifically, how to explain quantitatively the measured upper bounds in Figure 3.4 for $(x+\delta x)^{c}$, and Figure 3.4 for $\sin(x+\delta x)$?
@@
  • How to extend variance arithmetic to other input uncertainty distribution?
Traditional numerical approaches focus primarily on the result values, and they need to be reexamined or even reinvented in variance arithmetic. Most traditional numerical algorithms try to choose one of the best paths, while variance arithmetic rejects conceptually all path-dependent calculations. How to reconcile variance arithmetic and traditional numerical approaches could be a big challenge.
The calculation of the result mean and uncertainty contains a lot of sums of constant terms such as Formula (2.3) and (2.3), so that it is a excellent candidate for parallel processing. Because variance arithmetic has neither execution freedoms nor the associated dependency problems, it can be implemented either in software using GPU or directly in hardware.
In variance arithmetic, it is generally an order-of-magnitude more complicated to calculate the result uncertainties than the result values, such as Formula (6.19) for Taylor expansion of matrix inversion. However, modern software programs for analytic calculation such as SymPy can be a great help. Perhaps it is time to move from numerical programming to analytic programming for analytic problems.
@@ 10.8 Variable $\sigma$ Variance Arithmetic
Variance arithmetic rejects a calculation if its distributional pole or distributional zero is within the input ranges. But there are many use cases in which such intrusion cannot be avoided, so variance arithmetic of smaller $\sigma$ may be needed. It is even possible in another implementation of statistical Taylor expansion, $\sigma$ is part of a variance representation, so that a calculation closer to a distributional zero can proceed but at the expense of reducing $\sigma$. The result $\sigma$ for a calculation of imprecise inputs each with different $\sigma$ is found statistically through the corresponding $\epsilon$. Such approaches need theoretical foundation in both mathematics and statistics.